40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1421     Entry time: Tue Mar 24 13:10:07 2009
Author: Alberto 
Type: Configuration 
Category: PSL 
Subject: new mirror installed on the AP table 

New flipping mirror installed on the AP table on the beam path to the REFL199 PD.

If you're missing the double demod signal, please check that it is actually down.

ELOG V3.1.3-